40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 14973     Entry time: Wed Oct 16 11:42:17 2019     In reply to: 14962     Reply to this: 14974
Author: gautam 
Type: Update 
Category: LSC 
Subject: Poor separation of PRCL/MICH in 3f signals 

Summary:

There is poor separation of the PRCL and MICH length error signals as sensed in the 3f photodiodes. I don't know why this is so - one possibility is that the MICH-->PRM matrix element in the LSC output matrix needs to be tuned to minimize the MICH -->PRCL coupling.

Details:

Over the last few days, I've been trying to make the 3f locking of the PRMI more reliable. Turns out that while I was able to lock the PRMI on 3f error signals, it was just a fluke. So I set about trying to be more systematic. Here are the steps I followed:

  1. Lock the PRMI (i.e. ETMs misaligned) using REFL11 for PRCL, AS55 for MICH.
    • This is the so-called 1f scheme.
    • The servo signs are chosen such that the carrier field is resonant in the PRC.
    • Run the dither alignment to maximize POPDC, minimize ASDC. This is the main purpose of locking in this config.
    • Measure some loop TFs, make sure the servo gains are giving us ~100 Hz UGF on these loops.
  2. Change the sign of the servo loops to make the sidebands resonant in the PRC.
    • The error signals are still sourced from the 1f photodiodes.
    • Measure loop TFs, and also the TF between the 1f and 3f error signals. 
    • This allowed me to determine how the servo gains (and signs) that would be appropriate when using the 3f signals in place of the 1f.
    • Determine the offsets in the 3f error signals when the PRMI is locked on 1f error signals. This allows me to set the error point offsets for the PRCL_B and MICH_B paths, which are what is used for the 3f locking.
  3. Change the error signals from 1f to 3f. 
    • After much trial and error, I finally managed to get a stable (>10 mins) lock going.
    • Measured some loop TFs.
    • Turned on the notch filters in the control filter bank at the sensMat oscillator frequencies, and ran some sensing lines.

Attachment #1 is the result. I don't know what is the reason for such poor separation of the MICH and PRCL error signals in REFL165. The situation seems very different from when I had the DRMI locked in Nov last year.

After this exercise, I tried for some hours to get the 3f PRMI locking going with the arm cavities held off resonance under ALS control, but had no success. The angular motion of the PRC isn't helping, but I feel this shouldn't be a show stopper.

Attachment 1: sensMat.pdf  47 kB  Uploaded Wed Oct 16 12:55:51 2019  | Hide | Hide all
sensMat.pdf
ELOG V3.1.3-